simplify
4a2b/16ab2

Found 2 solutions by stanbon, haileytucki:
Answer by stanbon(75887)

simplify
4a^2b/16ab^2
-------------
= (4/16)(a^2/a)(b/b^2)
---
= (1/4)(a/1)(1/b)
---
= 1/(4b)
============
Cheers,
Stan H.

Answer by haileytucki(390)

(4a^(2)b)/(16ab^(2))
Reduce the expression (4a^(2)b)/(16ab^(2)) by removing a factor of 4ab from the numerator and denominator.
(a)/(4b)